OpenStudy (anonymous):

if a violin string Is 12inches long and vibrates 440 times per second what is the length shortened to 660 times per second

ganeshie8 (ganeshie8):

hint : \( \large f \propto \frac{1}{l}\)

ganeshie8 (ganeshie8):

\(f\) : frequency of vibration \(l\) : length of string

OpenStudy (anonymous):

so is it 440x1/12 ?

ganeshie8 (ganeshie8):

\(\large \frac{f_1}{f_2} = \frac{l_2}{l_1}\) \(\large l_1 \times \frac{f_1}{f_2} = l_2\)

ganeshie8 (ganeshie8):

\(\large 12 \times \frac{440}{660} = l_2\)

ganeshie8 (ganeshie8):

simplify completely

OpenStudy (anonymous):

0.061

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

12*440= 5280 / 600 = 8.8 ?

ganeshie8 (ganeshie8):

its **660

ganeshie8 (ganeshie8):

you need to divide by 660, not 600

OpenStudy (anonymous):

oh sorry Its very late and I saw it wrong

OpenStudy (anonymous):

its 8

ganeshie8 (ganeshie8):

8 is \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

and I have just one more problem

OpenStudy (anonymous):

do I have to repost a question or can I ask you it here

ganeshie8 (ganeshie8):

ask it here

OpenStudy (anonymous):

find equation for circle and line

OpenStudy (anonymous):

im trying to draw it it wont let me

ganeshie8 (ganeshie8):

there is a problem wid drawing tool from yesterday for some users

ganeshie8 (ganeshie8):

come here http://www.twiddla.com/1462359

OpenStudy (anonymous):